OpenStudy (coconut):

If f(x) = 5x – 4 and g(x) = 6x + 3, what is f(g(x))?

Parth (parthkohli):

= (5x -4)(6x + 3) = 5x(6x + 3) + -4(6x + 3) = 30x + 15x - 24x - 12 = 21x - 12

OpenStudy (coconut):

my options are 30x – 21 30x – 1 30x + 11 30x + 12

OpenStudy (anonymous):

30x + 12

Parth (parthkohli):

@Rohangrr can you please explain what went wrong in that? :(

Parth (parthkohli):

Oh yeah 5x * 6x = 30x2

OpenStudy (coconut):

can you show your work on how you got 30x+12

OpenStudy (anonymous):

see parth put on the formula!! then there is u r ans... i was also getting the same ans.. as u have but in the internet i got the correct formula and put the equation on it!! @ParthKohli

OpenStudy (coconut):

can someone show work on how they got that answer